OpenStudy (anonymous):

Simplify the radical expression: √14q * 2√4q PLS Help!

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

what are there please ? sqrt(14q) *2sqrt(4q) = ? this is right wrote hence ?

OpenStudy (anonymous):

so than this is wrote correct this will be equal = 2sqrt(14q*4q) = 2sqrt(2^2 *14*q^2) = 2q*2sqrt14 = 4q*sqrt14 hope this will help you good luck